OpenStudy (anonymous):

( (Sinx Cosy)+(Cosx Siny))/(Cosx Cosy) --------------------------------- ((Cosx Cosy)-(Sinx Siny))/(Cosx Cosy)

OpenStudy (jit4won):

cosx cos y will cancel each other

OpenStudy (jit4won):

sinxcosy +cosxsiny = sin(x+y) and cosx cosy -sinx siny = cos (x+y) so your final answer will be sin(x+y)/Cos(x+y)= tan (x+y)